A garden has the shape of a right triangle with one leg 7 meters longer than the other. The hypotenuse is 7 meters less than twice the length of the shorter leg. What is the length of the shorter leg?
Answer by lwsshak3(11628)

A garden has the shape of a right triangle with one leg 7 meters longer than the other. The hypotenuse is 7 meters less than twice the length of the shorter leg. What is the length of the shorter leg?
***
let x=length of shorter leg
x+7=length of longer leg
2x-7=length of hypotenuse
By the pythagorean theorem:
x^2+(x+7)^2=(2x-7)^2
x^2+x^2+14x+49=4x^2-28x+49
2x^2-42x=0
2x(x-21)=0
x=0 (reject)
or
x=21
x+7=28
2x-7=35
length of shorter leg=21 m
length of longer leg=28 m
length of hypotenuse=35 m